How many molecules are in 35.2 g Nh3?


1
Expert's answer
2021-05-03T07:32:20-0400

N=14, H=1, Molar mass = "14+1\times 3=17"

no. of moles "=35.2\/17=2.07moles"

no. of "NH_3 molecules=2.07\times6.02\times 10^{23}=1.246\times 10^{24} molecules"